write equation of line containing point (-2,4) and perpendicular to line 5x+2y=3
Answer by ad_alta(240)

In order to be perpendicular to the given line, the slope of our line must be the negative reciprocal of the given line. Thus, our slope should be 2/5. So far we have: y=(2/5)x+b. Plugging in the point to solve for 'b,' 4=(2/5)(-2)+b, so b=24/5 and our line is y=(2/5)*x+24/5.